OpenStudy (anonymous):

what is the product of 5/6 and 2/3? 10/12 5/9 7/9 15/12

OpenStudy (anonymous):

5/6 x 2/3 = (5x2)/(6x3)

OpenStudy (anonymous):

You just posted this problem and they showed you how to do it. You multiply the numerators first 5*2=10 Then you multiply the denominators 6*3=18 10/18, now you simplify this and get 5/9

OpenStudy (anonymous):

multiply the numerators then multiply the denominators